# gPC-based robustness analysis of neural systems through probabilistic recurrence metrics

### v1

@@ 3.2 Robustness at the cortical column level
The JR model (3), along with its extensions including interneurons and multiple cortical columns to better reproduce the activity of different brain regions, has been thoroughly analysed via bifurcation analysis, to investigate average inhibitory synaptic gain and time constant of mean postsynaptic potentials [subramaniyam2024sensitivity], as well as influence of input strength and dendritic time constant in a multi-input model [spiegler2010bifurcation], and input strength and interconnection gains in two coupled models [ahmadizadeh2018bifurcation]. None of these works, however, considers parametric uncertainty, which is our focus.
We study the robustness of the EEG-like signal $y(t)=x_{2}(t)-x_{3}(t)$ with respect to parametric uncertainty in the amplitude of excitatory and inhibitory postsynaptic potentials, $A$ and $B$, for the single-cortical-column JR model (3) from [jansen1995electroencephalogram]. Figure 7 shows the resulting PRP plots constructed, with the FG approach, in the parameter space $\mathcal{P}_{JR}=[A_{\min},A_{\max}]\times[B_{\min},B_{\max}]=[2,12]\times[15,31]$ also considered in [jansen1995electroencephalogram, Section 3.1], for different values of the input strength $p$ and of the connectivity constant $C$. Supplementary Figure S18 provides additional PRP plots for other values of $C$, while PRP plots obtained with the LE approach are in Supplementary Figure S17. Moreover, Supplementary Figure S2 demonstrates the three characteristic regimes of the JR model (3): alpha-wave (highest blob count, yellow regions in Figure 7, left), low-frequency wave (intermediate blob count, turquoise to green to orange regions in Figure 7, left), and no-wave (zero blob count in theory, white regions in Figure 7, left).
@@ 4 Conclusions
Focusing on dynamical systems in neuroscience, we have introduced a novel methodology for PRA aimed at quantifying how much probabilistic parametric uncertainty can be imposed on the nominal system parameter values, while ensuring that the mean system output preserves the qualitative recurrence pattern associated with the nominal dynamical regime. Assessing how much uncertainty the system can withstand while still preserving the nominal regime in expectation is crucial to analyse interpretable mathematical models in neuroscience under parametric uncertainty and identify the allowable uncertainty ranges for which a life-sustaining regime is preserved. We have employed the gPC approach to efficiently calculate mean neural signals through a simpler polynomial surrogate model. The gPC method has been recently employed to generate surrogate models for uncertain systems in neuroscience [ghori2023uncertainty, signorelli2024uncertainty, sutulovic2024efficient] and systems biology [paun2025comparison, sumser2025exploiting, paulson2019fast, hu2018generalized], and used for global sensitivity analysis of quantities of interest in specific mathematical models or for parameter estimation from data. Differently from other works, here we have relied on gPC surrogate models as a key step in our proposed general framework for the probabilistic robustness analysis of models in neuroscience. To detect loss of regime preservation from the gPC-computed mean neural time-series, we have developed an algorithm based on the identification and the persistence of geometric patterns in the associated recurrence plots; this has been necessary to overcome the impossibility of adopting traditional neuroscientific metrics used for deterministic signals, which cannot be computed or lose their interpretability in a probabilistic setting. From a computational perspective, to systematically identify geometric patterns, we have resorted to an automated count of the number of blobs, i.e., distinct regions with high values in recurrence plots; persistence of blob counts has been used as a metric for persistence of geometric patterns, and hence robust regime preservation. Finally, we have proposed two different approaches to investigate probabilistic robustness to parametric uncertainty within a whole parameter space of interest through probabilistic regime preservation (PRP) plots, which convey information both on the dynamical regimes that are preserved and on the level of tolerable parametric uncertainty affecting various nominal parameter values.
@@
The PRP plots for the membrane potential in the single-neuron HR model (2) reveal that plateau bursting is consistently more robust to probabilistic parametric uncertainty than square-wave bursting. Our novel methodology also provides an interesting insight: preservation of recurrence characteristics for mean state trajectories is unlikely in chaotic regions of a parameter space, as averaging completely different signals (arising even from very similar parameter values) leads to loss of any distinct geometric signature. The work [perinelli2020chasing] focuses on detection of chaos in human EEG brain signals and finds no evidence of chaotic behaviour, and our results (Figure 6) suggest an explanation for this absence of evidence: since a single electrode measures the mean activity of several neurons, a chaotic regime would be very difficult to observe experimentally, even when some individual neurons exhibit a chaotic firing pattern as in [gu2013biological].
@@
For the EEG-like signal generated by the single-cortical-column JR model (3), we investigate the effect of uncertainty in the amplitude of excitatory and inhibitory postsynaptic potentials. PRP plots (Figure 7) show how the regimes and the associated probabilistic robustness properties change for different values of the input and of the connectivity constant of the cortical column elements, and provide precious insight into the choice of the most appropriate parameter regions to consider when modelling the behaviour of cortical columns in a brain area of interest, based not only on the expected dynamical regimes, as in classical bifurcation analysis and in [jansen1995electroencephalogram], but also on the observed robustness properties under parametric uncertainties, unavoidable in experimental settings.
@@
In addition to what we have shown in this work, PRP plots can be precious to identify the most sensitive direction in which the nominal values of uncertain parameters can vary so as to transition from a dynamical regime to another, similarly to the analysis in [goldman2001global]. The novel information provided by our methodology is an estimate of the robustness to parametric uncertainty of the regime to be reached, which is fundamental for the design of control actions that safely steer the neural activity towards a desired functioning. For instance, to safely design an anaesthetic action that drives brain regions to a more relaxed dynamical regime [kuhlmann2016neural], we would like to minimise the changes with respect to nominal system parameter values, to avoid altering normal brain activity, while being reasonably sure to reach an anaesthetised state that is robust to parametric uncertainty, so that the control action can be effective under perturbations.
@@
Future research directions include refinements of the proposed PRA framework that enrich the information content of PRP plots: higher-order moments of state trajectories can be easily retrieved with gPC [lefebvre2020moment] (see, e.g., (5), where a formula for the variance of the stochastic process in (1) is also provided) and may be used to gain additional information that would, for instance, help distinguish between chaotic and quiescence regimes in a parameter space.

### v2

@@ 3.2 Robustness at the cortical column level
The JR model (3), along with its extensions including interneurons and multiple cortical columns to better reproduce the activity of different brain regions, has been thoroughly analysed via bifurcation analysis, to investigate average inhibitory synaptic gain and time constant of mean postsynaptic potentials [67], as well as influence of input strength and dendritic time constant in a multi-input model [66], and input strength and interconnection gains in two coupled models [2]. None of these works, however, considers parametric uncertainty, which is our focus.
We study the robustness of the EEG-like signal $y(t)=x_{2}(t)-x_{3}(t)$ with respect to parametric uncertainty in the amplitude of excitatory and inhibitory postsynaptic potentials, $A$ and $B$, for the single-cortical-column JR model (3) from [30]. Figure 7 shows the resulting PRP plots constructed, with the FG approach, in the parameter space $\mathcal{P}_{JR}=[A_{\min},A_{\max}]\times[B_{\min},B_{\max}]=[2,12]\times[15,31]$ also considered in [30, Section 3.1], for different values of the input strength $p$ and of the connectivity constant $C$. Supplementary Figure S18 provides additional PRP plots for other values of $C$, while PRP plots obtained with the LE approach are in Supplementary Figure S17. Moreover, Supplementary Figure S2 demonstrates the three characteristic regimes of the JR model (3): alpha-wave (highest blob count, yellow regions in Figure 7, left), low-frequency wave (intermediate blob count, turquoise to green to orange regions in Figure 7, left), and no-wave (zero blob count in theory, white regions in Figure 7, left).
@@ 4 Conclusions
Focusing on dynamical systems in neuroscience, we have introduced a novel methodology for PRA aimed at quantifying how much probabilistic parametric uncertainty can be imposed on the nominal system parameter values, while ensuring that the mean system output preserves the qualitative recurrence pattern associated with the nominal dynamical regime. Assessing how much uncertainty the system can withstand while still preserving the nominal regime in expectation is crucial to analyse interpretable mathematical models in neuroscience under parametric uncertainty and identify the allowable uncertainty ranges for which a life-sustaining regime is preserved. We have employed the gPC approach to efficiently calculate mean neural signals through a simpler polynomial surrogate model. The gPC method has been recently employed to generate surrogate models for uncertain systems in neuroscience [19, 64, 69] and systems biology [53, 68, 52, 27], and used for global sensitivity analysis of quantities of interest in specific mathematical models or for parameter estimation from data. Differently from other works, here we have relied on gPC surrogate models as a key step in our proposed general framework for the probabilistic robustness analysis of models in neuroscience. To detect loss of regime preservation from the gPC-computed mean neural time-series, we have developed an algorithm based on the identification and the persistence of geometric patterns in the associated recurrence plots; this has been necessary to overcome the impossibility of adopting traditional neuroscientific metrics used for deterministic signals, which cannot be computed or lose their interpretability in a probabilistic setting. From a computational perspective, to systematically identify geometric patterns, we have resorted to an automated count of the number of blobs, i.e., distinct regions with high values in recurrence plots; persistence of blob counts has been used as a metric for persistence of geometric patterns, and hence robust regime preservation. Finally, we have proposed two different approaches to investigate probabilistic robustness to parametric uncertainty within a whole parameter space of interest through probabilistic regime preservation (PRP) plots, which convey information both on the dynamical regimes that are preserved and on the level of tolerable parametric uncertainty affecting various nominal parameter values.
@@
The PRP plots for the membrane potential in the single-neuron HR model (2) reveal that plateau bursting is consistently more robust to probabilistic parametric uncertainty than square-wave bursting. Our novel methodology also provides an interesting insight: preservation of recurrence characteristics for mean state trajectories is unlikely in chaotic regions of a parameter space, as averaging completely different signals (arising even from very similar parameter values) leads to loss of any distinct geometric signature. The work [55] focuses on detection of chaos in human EEG brain signals and finds no evidence of chaotic behaviour, and our results (Figure 6) suggest an explanation for this absence of evidence: since a single electrode measures the mean activity of several neurons, a chaotic regime would be very difficult to observe experimentally, even when some individual neurons exhibit a chaotic firing pattern as in [23].
@@
For the EEG-like signal generated by the single-cortical-column JR model (3), we investigate the effect of uncertainty in the amplitude of excitatory and inhibitory postsynaptic potentials. PRP plots (Figure 7) show how the regimes and the associated probabilistic robustness properties change for different values of the input and of the connectivity constant of the cortical column elements, and provide precious insight into the choice of the most appropriate parameter regions to consider when modelling the behaviour of cortical columns in a brain area of interest, based not only on the expected dynamical regimes, as in classical bifurcation analysis and in [30], but also on the observed robustness properties under parametric uncertainties, unavoidable in experimental settings.
@@
In addition to what we have shown in this work, PRP plots can be precious to identify the most sensitive direction in which the nominal values of uncertain parameters can vary so as to transition from a dynamical regime to another, similarly to the analysis in [21]. The novel information provided by our methodology is an estimate of the robustness to parametric uncertainty of the regime to be reached, which is fundamental for the design of control actions that safely steer the neural activity towards a desired functioning. For instance, to safely design an anaesthetic action that drives brain regions to a more relaxed dynamical regime [38], we would like to minimise the changes with respect to nominal system parameter values, to avoid altering normal brain activity, while being reasonably sure to reach an anaesthetised state that is robust to parametric uncertainty, so that the control action can be effective under perturbations.
@@
Future research directions include refinements of the proposed PRA framework that enrich the information content of PRP plots: higher-order moments of state trajectories can be easily retrieved with gPC [39] (see, e.g., (5), where a formula for the variance of the stochastic process in (1) is also provided) and may be used to gain additional information that would, for instance, help distinguish between chaotic and quiescence regimes in a parameter space.